OpenStudy (anonymous):

Write a system of two equations using x and y to represent the following problem. A chemist is mixing two solutions together. The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?

OpenStudy (anonymous):

Hmm.. lets start out by seeing if you have any idea of where the X and Y might go

OpenStudy (anonymous):

Anything?

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

x+y=100 0.5x+0.36y=1 ??

OpenStudy (anonymous):

NICE!.. that was great,

OpenStudy (anonymous):

what do i do now?

OpenStudy (anonymous):

Add

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Do 0.50x +0.36y

OpenStudy (anonymous):

x+y=100 + 0.5x+0.36y=1 ---------------

OpenStudy (mathstudent55):

Let x = amount of 24% solution Let y = amount of 50% solution Total amount of solution made is: x + y = 100 Amount of solvent in solution: 0.24x + 0.50y = 0.36(100) 24x + 50y = 3600 12x + 25y = 1800 System of equations is: x + y = 100 12x + 25y = 1800

OpenStudy (anonymous):

someone help?!

OpenStudy (mathstudent55):

Solve the first equation for x: x = 100 - y Now substitute that into the second equation: 12(100 - y) + 25y = 1800 1200 - 12y + 25y = 1800 1200 + 13y = 1800 13y = 600 y = 600/13 y = 46.15 (approximately) x + y = 100 x + 600/13 = 100 x = 100 - 600/13 x = 1300/13 - 600/13 x = 700/13 x = 53.85 (approximately)

OpenStudy (anonymous):

put the amount of 24% solution as \(x\) gallons, then since the total is 100 gallons the amount of 50% solution must be \(100-x\) you want the ending amount to be 36% of 100 which is 36 set \[.24x+.5(100-x)=36\] and solve easier to multiply by 100 and solve \[24x+50(100-x)=3600\]

OpenStudy (anonymous):

\[24x+5000-50x=3600\] \[-26x=-1400\] \[x=\frac{1400}{26}\]

OpenStudy (mathmate):

You can do a mental calculation by subtracting 24 from 50 to get 26, which is the total of the two portions, in the ratios, 50-36=14. and 36-24=12 The volumes are therefore 100*14/26 and 100*12/26 (for x and y respectively).